OpenStudy (crashonce):

The HCF and LCM of three expressions are a^3bc^2 and a^5b^3c^4 respectively. if two of them are a^3bc^3 and a^4b^2c^4 then the third is?

OpenStudy (anonymous):

sorry what math is this?

OpenStudy (anonymous):

if its in a math book then you can look it up on Slader.com

OpenStudy (crashonce):

i want to know how to do it

OpenStudy (mathmath333):

its very tricky

OpenStudy (mathmath333):

their is is formula for when \(a\) and \(b\) are two positive integers then \(a\times b=HCF\times LCM\) but not that case if their are three if \(a\) ,\(b\) and \(c\) are three positive integers then \(a\times b\times c\neq HCF\times LCM\)

OpenStudy (crashonce):

no offense but that doesnt really help

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{x\times y\times z=\dfrac{LCM(x,y,z)\times HCF(x,y)\times HCF(y,z)\times HCF(z,x)}{HCF(x,y,z)}}\end{align}\)

OpenStudy (mathmath333):

is the answer or solution given

OpenStudy (loser66):

to me, it is \(a^5 b^3 c^2\)

OpenStudy (loser66):

if the first number is \(N_1=a^3bc^3\) and the second one is \(N_2=a^4b^2c^4\), then \(N_3=a^mb^nc^l\)

OpenStudy (loser66):

for LCM \(a^5b^3c^4\), it forces m=5, n=3 since the first term and the second term have the exponent of a and b <5

OpenStudy (loser66):

for HCF \(a^3bc^2\), it forces l =2 Hence the third one is \(a^5b^3c^2\)

OpenStudy (loser66):

*the exponent of a for \(N_1,N_2<5\) in LCM, that forces the exponent of a in \(N_3\) is 5

OpenStudy (loser66):

the exponent of b for \(N_1,N_2<3\) in LCM, that forces the exponent of b in \(N_3\) is 3

OpenStudy (loser66):

and the exponent of c relies on HCF which is 2

OpenStudy (mathmath333):

wait

OpenStudy (loser66):

yes, sir. :)

OpenStudy (loser66):

I don't know, but if c^4 is the answer on the third one, then the c in gcd is not c^2, it must be c^3

OpenStudy (mathmath333):

its not \(\huge c^4\) its is giving \(\huge z=ab\)

OpenStudy (loser66):